OpenStudy (anonymous):

If f(x) = x² + 1 and g(x) = 3x + 1, find f(2) + g(3).

OpenStudy (anonymous):

Plug a 2 into the f(x) function and a 3 into the g(x) function and add the two results together

OpenStudy (anonymous):

so 3(3)+1 and (2)sq+1?

OpenStudy (anonymous):

\[f(2)=2^2+1=4+1=5\] \[g(3)=3(3)+1=9+1=10\]